If a car at rest accelerates uniformly to a speed of 144 km/h in 20 s, it covers a distance of


        (a)    2880m      


(b)    1440m      


        (c)    400m       


(d)    20m


Grade:10

1 Answers

Jitender Pal
10 years ago

(c)
Initial velocity of car (u) = 0
Final velocity of car (v) = 144 Km/hr
= 40 m/s
Time taken = 20 s
We know that, v = u + at
40 = a × 20 => a = 2 m/s?
Also, v2 - u2 = 2as
    s=(v^2  –u^2)/2a
 s=((40)^2- (0)^2)/(2×2)=  1600/4=400 m.

(c)

Initial velocity of car (u) = 0

Final velocity of car (v) = 144 Km/hr

= 40 m/s

Time taken = 20 s

We know that, v = u + at

40 = a × 20 Þ a = 2 m/s?

Also, v2 - u2 = 2as

Þ    

Þ